OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

90 88x89x90=267

OpenStudy (unklerhaukus):

let the first number be \(n\) the second number is \(n+1\) {because the integers are consecutive {similarly } the third number is \(n+2\) summing \[(n)+(n+1)+(n+2)=267\] solving for \(n\) \[3n+3=267\] \[3n=264\] \[n=264/3=88\] so the three consecutive integers are \(n=88\) \(n+1=88+1=89\) \(n+2=88+2=90\)

OpenStudy (anonymous):

what is the largest one ? @UnkleRhaukus

OpenStudy (unklerhaukus):

are you asking which is larger 88, 89 or 90 ?

OpenStudy (anonymous):

largest integer? @UnkleRhaukus

OpenStudy (unklerhaukus):